trish30:

Identify the range of the exponential function y = {\left( {6.3} \right)^x} y = ( 6.3 ) x

Vocaloid:

can you try typing it out a little clearer? is it y = 6.3^x?

trish30:

sorry @Vocaloid you're right it is y=6.3^x

Vocaloid:

there are no restrictions on the domain, so we can examine the end behavior of the function if we plug in a very large negative value for x, the function approaches 0 but does not reach it if we plug in a very large positive value for x, the function approaches infinity

Vocaloid:

so your range should be (0,infinity)

trish30:

ok thanks but do you know how @Vocaloid

Vocaloid:

what do you mean?

Vocaloid:



Vocaloid:

I know because of prior knowledge/experience, but if you are uncertain, you can try plugging in values to see what you get 6.3^(-10000) = a very very small value (0.0000000000etcsomething) 6.3^(100000) = a very large value

Vocaloid:

it never reaches 0 so the range starts at (0) not inclusive

trish30:

@Vocaloid ok thanks